# Federated Time Series Generation on Feature and Temporally Misaligned Data

### v1

@@ 20      Finetune $\mathcal{D}$ on updated $\mathbf{X}^{\text{pub}}$
After this initial training, the coordinator distributes the trained global distiller model $\mathcal{D}$ to all participating clients. Each client $i$ then utilizes $\mathcal{D}$ to impute their missing common features. Since clients may have missing values in $\mathbf{X}_{\text{comm}}^{i}$, they input their data along with the corresponding mask $\mathbf{M}_{\text{comm}}^{i}$ to the distiller model, which will perform conditional generation to iteratively refine the imputed data by sampling from the conditional distribution guided by the observed data, shown in Equation 23. The imputation process follows $\hat{\mathbf{X}}_{\text{comm}}^{i}=\mathcal{D}(\mathbf{X}_{\text{comm}}^{i},\mathbf{M}_{\text{comm}}^{i})$, where $\mathcal{D}$ reconstructs only the missing values, indicated by $\mathbf{M}_{\text{comm}}^{i}=0$. Similarly, the local imputer imputes missing values in $\mathbf{X}_{\text{ex}}^{i}$ by inputting their data along with the corresponding mask $\mathbf{M}_{\text{ex}}^{i}$ to the imputer model via $\hat{\mathbf{X}}_{\text{ex}}^{i}=\mathcal{U}(\mathbf{X}_{\text{ex}}^{i},\mathbf{M}_{\text{ex}}^{i})$. The imputed common features $\hat{\mathbf{X}}_{\text{comm}}^{i}$ are then combined with the available exclusive features $\hat{\mathbf{X}}_{\text{ex}}^{i}$ to form the training data $\mathbf{X}^{i}_{\text{train}}=\hat{\mathbf{X}}_{\text{comm}}^{i}\cup\hat{\mathbf{X}}_{\text{ex}}^{i}$ for the local imputer. Meanwhile, each client trains their local imputer model $\mathcal{U}^{i}$ using $\mathbf{X}^{i}_{\text{train}}$ as the ground truth. Since the imputed common features $\hat{\mathbf{X}}_{\text{comm}}^{i}$ are fully known (as they are outputs from the pre-trained and fine-tuned $\mathcal{D}$), they are entirely used as ground truth for training $\mathcal{U}^{i}$, regardless of the original mask $\mathbf{M}_{\text{comm}}^{i}$. For the exclusive features, only the observed entries indicated by the mask $\mathbf{M}_{\text{ex}}^{i}$ are used as ground truth since the quality of the imputer’s generated data during training is not sufficient to be used as ground truth. We define the loss mask as $\mathbf{M}^{i}_{\text{loss}}=\mathbf{1}_{\text{comm}}^{i}\cup\mathbf{M}_{\text{ex}}^{i}$, where $\mathbf{1}_{\text{comm}}^{i}$ is a matrix of ones corresponding to the common features of client $i$. This loss mask ensures that the reconstruction loss is computed over all entries of the imputed common features and the observed entries of the exclusive features. The training loss for the imputer $\mathcal{U}^{i}$ can be defined as follows:
@@
$\text{where}\quad\mathcal{L}_{\text{time}}^{i}=\mathbb{E}_{(j,k,t)\,|\,\mathbf{M}^{i}_{\text{loss}_{j,k}}=1}\left[\left\|\mathbf{X}^{i}_{\text{train}_{j,k}}-\tilde{\mathbf{X}}^{i}_{\text{train}_{j,k}}(\mathbf{X}^{i}_{\text{train}_{j,k,t}},t;\theta)\right\|^{2}\right]$ (5)
$\text{and}\quad\mathcal{L}_{\text{freq}}^{i}=\mathbb{E}_{(j,k,t)\,|\,\mathbf{M}^{i}_{\text{loss}_{j,k}}=1}\left[\left\|\text{FFT}(\mathbf{X}^{i}_{\text{train}_{j,k}})-\text{FFT}\left(\tilde{\mathbf{X}}^{i}_{\text{train}_{j,k}}(\mathbf{X}^{i}_{\text{train}_{j,k,t}},t;\theta)\right)\right\|^{2}\right],$ (6)
where $\mathbf{X}^{i}_{\text{train}_{j,k}}$ is the $(j,k)$-th entry of $\mathbf{X}^{i}_{\text{train}}$, $\tilde{\mathbf{X}}^{i}_{\text{train}_{j,k}}$ is the denoised estimate from $\mathcal{U}$. After training, each client uses the trained imputer $\mathcal{U}^{i}$ to generate a synthetic dataset through unconditional synthesis, which includes both the common features $\hat{\mathbf{X}}_{\text{comm}}^{i}$ and the exclusive features $\hat{\mathbf{X}}_{\text{ex}}^{i}$. Starting from Gaussian noise, the imputer generates samples $\hat{\mathbf{X}}^{i}=\mathcal{U}^{i}(\mathbf{z}),\mathbf{z}\sim\mathcal{N}(0,\mathbf{I})$, that capture the distribution of both common and exclusive features.
@@ 4 Experiments
To assess the quality of synthetic data, we consider four real-world datasets and one simulated dataset with different properties, such as the number of features, correlation, periodicity, and noise levels. Each dataset is preprocessed using a sliding window technique (Yoon et al., 2019) to segment the data into sequences of length 24 to capture meaningful temporal dependencies while keeping the computational cost manageable. Stocks (Yoon et al., 2019) is the daily historical Google stock data from 2004 to 2019 with highly correlated features. ETTh (Zhou et al., 2021) recorded the electricity transformers hourly between July 2016 and July 2018, including load and oil temperature data that consists of 7 features. Energy (Candanedo, 2017) from UCI appliances energy prediction dataset with 10-minute intervals for about 4.5 months. fMRI (Smith et al., 2011) is a realistic simulation of brain activity time series with 50 features. MuJoCo (Tunyasuvunakool et al., 2020) is a physics-based simulation time series containing 14 features. We show the statistics of all datasets in Appendix D.2.
@@
We compare FedTDD against approaches show in 3(a), 3(b), 3(c) and 3(d). For the Centralized* training, we aggregate all data from individual clients, including public data, into a single location, where a global model is trained using the combined dataset, and this will be trained with all available features in the dataset and without missing values. While Centralized uses the same training procedure as Centralized*, it is, however, trained on a combined dataset with missing values and corresponding features available from each client plus the public data. To deal with differing features across clients, we create the combined dataset consisting of the total number of features in the particular benchmark dataset and zero-fill any remaining features to ensure uniformity. On the other hand, Local training involves training a separate model for each client using only their local data, without any communication or data aggregation. This approach has to be done to verify that FedTDD can perform relatively better than train locally. Finally, the Pre-trained approach leverages a model trained on a public dataset and uses it to impute the common features in local data from each client. Again, there is no data aggregation for this approach. In comparison, FedTDD integrates the Pre-trained approach and applies data aggregation to it. We utilized a SOTA diffusion-based multivariate time series generative model, Diffusion-TS (Yuan & Qiao, 2024), as the backbone for these baselines and FedTDD. Alternatively, any other time series generative model can be adopted in these approaches in a plug-and-play manner.
@@
We quantitatively assess the quality of the generated synthetic data using four key metrics (see Appendix D.3 for more details). Context-Fréchet Inception Distance (Context-FID) score (Jeha et al., 2022) evaluates the similarity between the distribution of real and synthetic time series data by computing the Fréchet distance. Correlational score (Liao et al., 2020) measures the correlation between the features of multivariate time series in the synthetic data compared to its real data. Discriminative score (Yoon et al., 2019) measures the realism of the synthetic data by training a binary classifier to distinguish between real and synthetic data. Predictive score (Yoon et al., 2019) evaluates the utility of the synthetic data by training a sequence-to-sequence model on the synthetic data and measuring its performance on real data. All evaluation metrics are computed based on the respective features of the individual clients and then averaged over five trials, followed by calculating the overall average across the number of clients. The quality of synthetic data is considered the “best” when all metrics approach 0, meaning lower values indicate better quality.
@@ 4.1 Time Series Generation
Context-FID Centralized* 0.682 +/- 0.106 0.281 +/- 0.040 0.782 +/- 0.138 0.533 +/- 0.082 1.737 +/- 0.125
Centralized 3.548 +/- 0.990 8.870 +/- 2.295 10.00 +/- 2.814 9.343 +/- 2.808 13.56 +/- 3.357
Local 1.648 +/- 0.229 1.313 +/- 0.188 0.751 +/- 0.121 1.179 +/- 0.179 1.694 +/- 0.153
Pre-trained 1.047 +/- 0.169 0.326 +/- 0.040 0.617 +/- 0.090 0.412 +/- 0.054 1.411 +/- 0.102
FedTDD 0.675 +/- 0.087 0.271 +/- 0.038 0.529 +/- 0.068 0.376 +/- 0.056 1.459 +/- 0.099
Correlational Centralized* 0.061 +/- 0.043 0.253 +/- 0.094 1.989 +/- 0.247 5.231 +/- 1.294 7.900 +/- 0.384
Centralized 0.769 +/- 0.336 0.340 +/- 0.097 2.230 +/- 0.518 5.681 +/- 0.634 18.07 +/- 2.311
Local 0.156 +/- 0.120 0.239 +/- 0.079 1.298 +/- 0.260 3.447 +/- 0.838 5.992 +/- 0.383
Pre-trained 0.077 +/- 0.052 0.165 +/- 0.074 1.323 +/- 0.171 2.821 +/- 0.651 6.049 +/- 0.349
FedTDD 0.058 +/- 0.050 0.161 +/- 0.064 1.296 +/- 0.215 2.800 +/- 0.686 6.017 +/- 0.364
Discriminative Centralized* 0.136 +/- 0.091 0.199 +/- 0.061 0.297 +/- 0.108 0.230 +/- 0.080 0.422 +/- 0.074
Centralized 0.476 +/- 0.042 0.475 +/- 0.017 0.474 +/- 0.024 0.496 +/- 0.006 0.477 +/- 0.030
Local 0.340 +/- 0.153 0.298 +/- 0.060 0.200 +/- 0.092 0.329 +/- 0.087 0.397 +/- 0.061
Pre-trained 0.175 +/- 0.117 0.115 +/- 0.060 0.208 +/- 0.068 0.141 +/- 0.068 0.419 +/- 0.051
FedTDD 0.185 +/- 0.105 0.106 +/- 0.061 0.153 +/- 0.120 0.153 +/- 0.072 0.414 +/- 0.051
Predictive Centralized* 0.040 +/- 0.000 0.127 +/- 0.003 0.112 +/- 0.015 0.292 +/- 0.009 0.137 +/- 0.004
Centralized 0.047 +/- 0.012 0.223 +/- 0.020 0.165 +/- 0.060 0.427 +/- 0.053 0.233 +/- 0.051
Local 0.043 +/- 0.003 0.118 +/- 0.011 0.048 +/- 0.006 0.204 +/- 0.012 0.135 +/- 0.006
Pre-trained 0.046 +/- 0.001 0.104 +/- 0.004 0.052 +/- 0.004 0.177 +/- 0.005 0.133 +/- 0.006
FedTDD 0.041 +/- 0.001 0.101 +/- 0.004 0.048 +/- 0.004 0.175 +/- 0.006 0.133 +/- 0.004
@@ 4.2 Ablation Study
Context-FID Centralized* 0.682 +/- 0.106 0.281 +/- 0.040 0.782 +/- 0.138 0.533 +/- 0.082 1.737 +/- 0.125
Centralized 3.733 +/- 0.959 11.54 +/- 3.894 14.68 +/- 4.263 13.17 +/- 3.035 15.34 +/- 4.789
Local 1.982 +/- 0.234 0.824 +/- 0.105 0.660 +/- 0.100 0.844 +/- 0.127 1.220 +/- 0.098
Pre-trained 0.738 +/- 0.142 0.316 +/- 0.032 0.547 +/- 0.099 0.381 +/- 0.066 1.178 +/- 0.104
FedTDD 0.680 +/- 0.123 0.267 +/- 0.036 0.510 +/- 0.072 0.331 +/- 0.051 1.196 +/- 0.098
Correlational Centralized* 0.061 +/- 0.043 0.253 +/- 0.094 1.989 +/- 0.247 5.231 +/- 1.294 7.900 +/- 0.384
Centralized 0.697 +/- 0.168 0.523 +/- 0.095 2.317 +/- 0.597 5.781 +/- 0.924 31.35 +/- 4.923
Local 0.091 +/- 0.052 0.167 +/- 0.057 1.079 +/- 0.196 1.984 +/- 0.594 4.929 +/- 0.395
Pre-trained 0.028 +/- 0.027 0.132 +/- 0.054 1.115 +/- 0.233 1.795 +/- 0.577 5.033 +/- 0.323
FedTDD 0.025 +/- 0.022 0.137 +/- 0.064 1.060 +/- 0.209 1.737 +/- 0.282 5.005 +/- 0.317
Discriminative Centralized* 0.136 +/- 0.091 0.199 +/- 0.061 0.297 +/- 0.108 0.230 +/- 0.080 0.422 +/- 0.074
Centralized 0.475 +/- 0.041 0.469 +/- 0.020 0.479 +/- 0.026 0.494 +/- 0.010 0.484 +/- 0.023
Local 0.300 +/- 0.116 0.208 +/- 0.070 0.190 +/- 0.088 0.241 +/- 0.071 0.398 +/- 0.058
Pre-trained 0.119 +/- 0.088 0.116 +/- 0.067 0.163 +/- 0.088 0.130 +/- 0.058 0.418 +/- 0.050
FedTDD 0.112 +/- 0.097 0.107 +/- 0.078 0.157 +/- 0.104 0.120 +/- 0.067 0.412 +/- 0.057
Predictive Centralized* 0.040 +/- 0.000 0.127 +/- 0.003 0.112 +/- 0.015 0.292 +/- 0.009 0.137 +/- 0.004
Centralized 0.168 +/- 0.025 0.196 +/- 0.027 0.198 +/- 0.049 0.314 +/- 0.052 0.223 +/- 0.029
Local 0.084 +/- 0.038 0.114 +/- 0.009 0.069 +/- 0.010 0.199 +/- 0.007 0.130 +/- 0.005
Pre-trained 0.028 +/- 0.007 0.108 +/- 0.004 0.063 +/- 0.007 0.190 +/- 0.005 0.132 +/- 0.005
FedTDD 0.028 +/- 0.005 0.107 +/- 0.005 0.062 +/- 0.006 0.186 +/- 0.004 0.130 +/- 0.005

### v2

@@ 20 Finetune $\mathcal{D}$ on updated $\mathbf{X}^{\text{pub}}$
After this initial training, the coordinator distributes the trained global distiller model $\mathcal{D}$ to all participating clients. Each client $i$ then utilizes $\mathcal{D}$ to impute their missing common features. Since clients may have missing values in $\mathbf{X}_{\text{comm}}^{i}$, they input their data along with the corresponding mask $\mathbf{M}_{\text{comm}}^{i}$ to the distiller model, which will perform conditional generation to iteratively refine the imputed data by sampling from the conditional distribution guided by the observed data, shown in Equation 17 in Appendix B.2. The imputation process follows $\hat{\mathbf{X}}_{\text{comm}}^{i}=\mathcal{D}(\mathbf{X}_{\text{comm}}^{i},\mathbf{M}_{\text{comm}}^{i})$, where $\mathcal{D}$ reconstructs only the missing values, indicated by $\mathbf{M}_{\text{comm}}^{i}=0$. Similarly, the local imputer imputes missing values in $\mathbf{X}_{\text{ex}}^{i}$ by inputting their data along with the corresponding mask $\mathbf{M}_{\text{ex}}^{i}$ to the imputer model via $\hat{\mathbf{X}}_{\text{ex}}^{i}=\mathcal{U}(\mathbf{X}_{\text{ex}}^{i},\mathbf{M}_{\text{ex}}^{i})$. The imputed common features $\hat{\mathbf{X}}_{\text{comm}}^{i}$ are then combined with the available exclusive features $\hat{\mathbf{X}}_{\text{ex}}^{i}$ to form the training data $\mathbf{X}^{i}_{\text{train}}=\hat{\mathbf{X}}_{\text{comm}}^{i}\cup\hat{\mathbf{X}}_{\text{ex}}^{i}$ for the local imputer. Meanwhile, each client trains their local imputer model $\mathcal{U}^{i}$ using $\mathbf{X}^{i}_{\text{train}}$ as the ground truth. Since the imputed common features $\hat{\mathbf{X}}_{\text{comm}}^{i}$ are fully known (as they are outputs from the pre-trained and fine-tuned $\mathcal{D}$), they are entirely used as ground truth for training $\mathcal{U}^{i}$, regardless of the original mask $\mathbf{M}_{\text{comm}}^{i}$. For the exclusive features, only the observed entries indicated by the mask $\mathbf{M}_{\text{ex}}^{i}$ are used as ground truth since the quality of the imputer’s generated data during training is not sufficient to be used as ground truth. We define the loss mask as $\mathbf{M}^{i}_{\text{loss}}=\mathbf{1}_{\text{comm}}^{i}\cup\mathbf{M}_{\text{ex}}^{i}$, where $\mathbf{1}_{\text{comm}}^{i}$ is a matrix of ones corresponding to the common features of client $i$. This loss mask ensures that the reconstruction loss is computed over all entries of the imputed common features and the observed entries of the exclusive features. The training loss for the imputer $\mathcal{U}^{i}$ can be defined as follows:
@@
$\text{where}\quad\mathcal{L}_{\text{time}}^{i}=\mathbb{E}_{(j,k,t)\,|\,\mathbf{M}^{i}_{\text{loss}_{j,k}}=1}\left[\left\|\mathbf{X}^{i}_{\text{train}_{j,k}}-\tilde{\mathbf{X}}^{i}_{\text{train}_{j,k}}(\mathbf{X}^{i}_{\text{train}_{j,k,t}},t;\theta)\right\|^{2}\right]$
$\text{and}\quad\mathcal{L}_{\text{freq}}^{i}=\mathbb{E}_{(j,k,t)\,|\,\mathbf{M}^{i}_{\text{loss}_{j,k}}=1}\left[\left\|\text{FFT}(\mathbf{X}^{i}_{\text{train}_{j,k}})-\text{FFT}\left(\tilde{\mathbf{X}}^{i}_{\text{train}_{j,k}}(\mathbf{X}^{i}_{\text{train}_{j,k,t}},t;\theta)\right)\right\|^{2}\right],$
[…]
@@ 4 Experiments
To assess the quality of synthetic data, we consider four real-world datasets and one simulated dataset with different properties, such as the number of features, correlation, periodicity, and noise levels. Each dataset is preprocessed using a sliding window technique [34] to segment the data into sequences of length 24 to capture meaningful temporal dependencies while keeping the computational cost manageable. Stocks is the daily historical Google stock data from 2004 to 2019 with highly correlated features. ETTh recorded the electricity transformers hourly between July 2016 and July 2018, including load and oil temperature data that consists of 7 features. Energy from UCI appliances energy prediction dataset with 10-minute intervals for about 4.5 months. fMRI is a realistic simulation of brain activity time series with 50 features. MuJoCo is a physics-based simulation time series containing 14 features. We show the statistics of all datasets in Appendix D.2.
@@
We compare FedTDD against approaches show in 3(a), 3(b), 3(c) and 3(d). For the Centralized* training, we aggregate all data from individual clients, including public data, into a single location, where a global model is trained using the combined dataset, and this will be trained with all available features in the dataset and without missing values. While Centralized uses the same training procedure as Centralized*, it is, however, trained on a combined dataset with missing values and corresponding features available from each client plus the public data. To deal with differing features across clients, we create the combined dataset consisting of the total number of features in the particular benchmark dataset and zero-fill any remaining features to ensure uniformity. On the other hand, Local training involves training a separate model for each client using only their local data, without any communication or data aggregation. This approach has to be done to verify that FedTDD can perform relatively better than train locally. Finally, the Pre-trained approach leverages a model trained on a public dataset and uses it to impute the common features in local data from each client. Again, there is no data aggregation for this approach. In comparison, FedTDD integrates the Pre-trained approach and applies data aggregation to it. We utilized a SOTA diffusion-based multivariate time series generative model, Diffusion-TS [35], as the backbone for these baselines and FedTDD. Alternatively, any other time series generative model can be adopted in these approaches in a plug-and-play manner.
@@
We quantitatively assess the quality of the generated synthetic data using four key metrics (see Appendix D.3 for more details). Context-Fréchet Inception Distance (Context-FID) score [10] evaluates the similarity between the distribution of real and synthetic time series data by computing the Fréchet distance. Correlational score [15] measures the correlation between the features of multivariate time series in the synthetic data compared to its real data. Discriminative score [34] measures the realism of the synthetic data by training a binary classifier to distinguish between real and synthetic data. Predictive score [34] evaluates the utility of the synthetic data by training a sequence-to-sequence model on the synthetic data and measuring its performance on real data. All evaluation metrics are computed based on the respective features of the individual clients and then averaged over five trials, followed by calculating the overall average across the number of clients. The quality of synthetic data is considered the “best” when all metrics approach 0, meaning lower values indicate better quality.
@@ 4.1 Time Series Generation
Context-FID Centralized* 0.682$\pm$0.106 0.281$\pm$0.040 0.782$\pm$0.138 0.533$\pm$0.082 1.737$\pm$0.125
Centralized 3.548$\pm$0.990 8.870$\pm$2.295 10.00$\pm$2.814 9.343$\pm$2.808 13.56$\pm$3.357
Local 1.648$\pm$0.229 1.313$\pm$0.188 0.751$\pm$0.121 1.179$\pm$0.179 1.694$\pm$0.153
Pre-trained 1.047$\pm$0.169 0.326$\pm$0.040 0.617$\pm$0.090 0.412$\pm$0.054 1.411$\pm$0.102
FedTDD 0.675$\pm$0.087 0.271$\pm$0.038 0.529$\pm$0.068 0.376$\pm$0.056 1.459$\pm$0.099
Correlational Centralized* 0.061$\pm$0.043 0.253$\pm$0.094 1.989$\pm$0.247 5.231$\pm$1.294 7.900$\pm$0.384
Centralized 0.769$\pm$0.336 0.340$\pm$0.097 2.230$\pm$0.518 5.681$\pm$0.634 18.07$\pm$2.311
Local 0.156$\pm$0.120 0.239$\pm$0.079 1.298$\pm$0.260 3.447$\pm$0.838 5.992$\pm$0.383
Pre-trained 0.077$\pm$0.052 0.165$\pm$0.074 1.323$\pm$0.171 2.821$\pm$0.651 6.049$\pm$0.349
FedTDD 0.058$\pm$0.050 0.161$\pm$0.064 1.296$\pm$0.215 2.800$\pm$0.686 6.017$\pm$0.364
Discriminative Centralized* 0.136$\pm$0.091 0.199$\pm$0.061 0.297$\pm$0.108 0.230$\pm$0.080 0.422$\pm$0.074
Centralized 0.476$\pm$0.042 0.475$\pm$0.017 0.474$\pm$0.024 0.496$\pm$0.006 0.477$\pm$0.030
Local 0.340$\pm$0.153 0.298$\pm$0.060 0.200$\pm$0.092 0.329$\pm$0.087 0.397$\pm$0.061
Pre-trained 0.175$\pm$0.117 0.115$\pm$0.060 0.208$\pm$0.068 0.141$\pm$0.068 0.419$\pm$0.051
FedTDD 0.185$\pm$0.105 0.106$\pm$0.061 0.153$\pm$0.120 0.153$\pm$0.072 0.414$\pm$0.051
Predictive Centralized* 0.040$\pm$0.000 0.127$\pm$0.003 0.112$\pm$0.015 0.292$\pm$0.009 0.137$\pm$0.004
Centralized 0.047$\pm$0.012 0.223$\pm$0.020 0.165$\pm$0.060 0.427$\pm$0.053 0.233$\pm$0.051
Local 0.043$\pm$0.003 0.118$\pm$0.011 0.048$\pm$0.006 0.204$\pm$0.012 0.135$\pm$0.006
Pre-trained 0.046$\pm$0.001 0.104$\pm$0.004 0.052$\pm$0.004 0.177$\pm$0.005 0.133$\pm$0.006
FedTDD 0.041$\pm$0.001 0.101$\pm$0.004 0.048$\pm$0.004 0.175$\pm$0.006 0.133$\pm$0.004
@@ 4.2 Ablation Study
Context-FID Centralized* 0.682$\pm$0.106 0.281$\pm$0.040 0.782$\pm$0.138 0.533$\pm$0.082 1.737$\pm$0.125
Centralized 3.733$\pm$0.959 11.54$\pm$3.894 14.68$\pm$4.263 13.17$\pm$3.035 15.34$\pm$4.789
Local 1.982$\pm$0.234 0.824$\pm$0.105 0.660$\pm$0.100 0.844$\pm$0.127 1.220$\pm$0.098
Pre-trained 0.738$\pm$0.142 0.316$\pm$0.032 0.547$\pm$0.099 0.381$\pm$0.066 1.178$\pm$0.104
FedTDD 0.680$\pm$0.123 0.267$\pm$0.036 0.510$\pm$0.072 0.331$\pm$0.051 1.196$\pm$0.098
Correlational Centralized* 0.061$\pm$0.043 0.253$\pm$0.094 1.989$\pm$0.247 5.231$\pm$1.294 7.900$\pm$0.384
Centralized 0.697$\pm$0.168 0.523$\pm$0.095 2.317$\pm$0.597 5.781$\pm$0.924 31.35$\pm$4.923
Local 0.091$\pm$0.052 0.167$\pm$0.057 1.079$\pm$0.196 1.984$\pm$0.594 4.929$\pm$0.395
Pre-trained 0.028$\pm$0.027 0.132$\pm$0.054 1.115$\pm$0.233 1.795$\pm$0.577 5.033$\pm$0.323
FedTDD 0.025$\pm$0.022 0.137$\pm$0.064 1.060$\pm$0.209 1.737$\pm$0.282 5.005$\pm$0.317
Discriminative Centralized* 0.136$\pm$0.091 0.199$\pm$0.061 0.297$\pm$0.108 0.230$\pm$0.080 0.422$\pm$0.074
Centralized 0.475$\pm$0.041 0.469$\pm$0.020 0.479$\pm$0.026 0.494$\pm$0.010 0.484$\pm$0.023
Local 0.300$\pm$0.116 0.208$\pm$0.070 0.190$\pm$0.088 0.241$\pm$0.071 0.398$\pm$0.058
Pre-trained 0.119$\pm$0.088 0.116$\pm$0.067 0.163$\pm$0.088 0.130$\pm$0.058 0.418$\pm$0.050
FedTDD 0.112$\pm$0.097 0.107$\pm$0.078 0.157$\pm$0.104 0.120$\pm$0.067 0.412$\pm$0.057
Predictive Centralized* 0.040$\pm$0.000 0.127$\pm$0.003 0.112$\pm$0.015 0.292$\pm$0.009 0.137$\pm$0.004
Centralized 0.168$\pm$0.025 0.196$\pm$0.027 0.198$\pm$0.049 0.314$\pm$0.052 0.223$\pm$0.029
Local 0.084$\pm$0.038 0.114$\pm$0.009 0.069$\pm$0.010 0.199$\pm$0.007 0.130$\pm$0.005
Pre-trained 0.028$\pm$0.007 0.108$\pm$0.004 0.063$\pm$0.007 0.190$\pm$0.005 0.132$\pm$0.005
FedTDD 0.028$\pm$0.005 0.107$\pm$0.005 0.062$\pm$0.006 0.186$\pm$0.004 0.130$\pm$0.005